OpenStudy (anonymous):

PLEASE HELP!!! Solve Complex Numbers √−112

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

Try splitting the \[\sqrt{-122}\] into 2 square roots, one involving \[\sqrt{-1}\] and one involving the square root of something else

OpenStudy (anonymous):

would the other one be \[\sqrt{122}\]

OpenStudy (anonymous):

There's no way to solve \[\sqrt{-122}\]

OpenStudy (anonymous):

although...

OpenStudy (anonymous):

it invovles immaginary numbers

OpenStudy (anonymous):

you can do \[\sqrt{-1}* \sqrt{122}\]

OpenStudy (anonymous):

\[\sqrt{-1} = i\]

OpenStudy (anonymous):

what's the sqrt of 122?

OpenStudy (anonymous):

You are correct, the other square root would be \[\sqrt{122}\] Try simplifying it now

OpenStudy (anonymous):

\[11.0453 i\] is your answer

OpenStudy (anonymous):

its the squareroot 112 not 122

OpenStudy (anonymous):

Whoops, my apologies. Can \[\sqrt{112}\] be simplified any further?

OpenStudy (anonymous):

oh then it would be 10.58300524 i

OpenStudy (anonymous):

its 4\[\sqrt{7}\]

OpenStudy (anonymous):

4 squareroot 7

OpenStudy (anonymous):

Looks good, don't forget the imaginary i :)

OpenStudy (anonymous):

do i multiply it by \[\sqrt{-1}\] or \[\sqrt{i}\]

OpenStudy (anonymous):

\[\sqrt{-1} = i\] so you'll want to multiply by i

OpenStudy (anonymous):

thank you